OpenStudy (anonymous):

solve (0,2pi] 2 cos x tan x +square3 tan x =0 :( help

OpenStudy (anonymous):

between both terms, tan x is common. So factor out tan x.

OpenStudy (anonymous):

\[\tan x (2 \cos x +\sqrt{3})\]

OpenStudy (anonymous):

Great. Now set each factor equal to 0.

OpenStudy (anonymous):

tanx=0 cosx = -\[\sqrt{3/2}\]

OpenStudy (anonymous):

- sqrt(3)/2

OpenStudy (anonymous):



OpenStudy (anonymous):

now find the angles x that satisfies each equation.

OpenStudy (anonymous):

cosx = \[5\pi/6 , 7\pi/6\]

OpenStudy (anonymous):

great! what about tan x = 0?

OpenStudy (anonymous):

tanx = cos/sin

OpenStudy (anonymous):

dont go that route

OpenStudy (anonymous):

you should know, that tan 0 = 0; tan 180 = 0, tan 360 = 0

OpenStudy (anonymous):

but you dont want to include 0 as the directions say (0,360]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Look at the tan x graph

OpenStudy (anonymous):



OpenStudy (anonymous):

you can see from the graph that tan 0 = 0; tan 180 (pi) = 0; tan 360 (2 pi) = 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So you have 4 solutions to the equation.

OpenStudy (anonymous):

5pi/6, 7pi/6 , 0

OpenStudy (anonymous):

No...we said that 0 is not to be included. You need to include pi and 2 pi. (thats where tan x = 0)

OpenStudy (anonymous):

pi and 2pi

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you understand why we are excluding 0 for this problem?

OpenStudy (anonymous):

i'm

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

thanks for you time :)